# GAD-Generative Learning for HD Map-Free Autonomous Driving

### v2

@@ 4.3 Closed-loop Experimental
Initially, we conducted an ablation study to assess the necessity of various samplers within a closed road environment. Subsequently, we identified the optimal sampler combination for further evaluation on open urban roadways. The outcomes of this ablation study are delineated in the appendix. Comprehensive real-world testing of these methodologies was conducted in densely populated urban settings, under the vigilant supervision of human safety drivers. The specific driving route utilized for our experiments is depicted in the appendix. Throughout the extensive public road testing, spanning over 30 kilometers, the model adeptly executed a diverse array of complex maneuvers, such as lane-following, merging, yielding to pedestrians. The results of these road tests are documented in Tab. 3
@@
It is noteworthy that methodologies exhibiting superior imitation metrics do not invariably demonstrate enhanced performance in closed-loop evaluations. This discrepancy may be attributable to the limitations of the imitation metric in adequately assessing the complexities of planning tasks. Nonetheless, the multi-modal imitation generator and GAN-based generator significantly enhance the system’s performance in both open-loop and closed-loop settings, primarily in terms of the planner’s intelligence and the continuity between consecutive frames. The efficiency and comfort metrics both show improvements with the integration of data-driven generators. Closed-loop results suggest that multimodality mitigates the brittleness inherent in imitation learning, while the lattice sampler further enhances the robustness of the entire system.
@@
To rigorously evaluate the efficacy of the proposed methodology as a substitute for off-the-shelf planners, we conducted a comparative analysis between the GAD and a conventional modularized, user-level, rule-based system, referred to as DPQP in Tab. 3. Notwithstanding, a discernible performance discrepancy persists in terms of system’s robustness and comfortableness. We posit that increase training data size could ameliorate this disparity.
@@ 5 Conclusions
We introduce a sophisticated DL pipeline from prediction to motion planning in HD map-free setting, augmented with an enhanced safety assurance. Our investigations have yielded substantial enhancements over prevailing data-driven methodologies within the complex urban environments of Shanghai. This paradigm enables the safety deployment of learned planners in real-world scenarios, offering interpretability and leveraging their capacity to evolve with accumulating data, thereby managing scenarios of greater complexity than those addressed by traditional rule-based systems. Future work may integrate generative learning techniques such as Generative Adversarial Networks (GAN), Conditional Variational Autoencoders (CVAE), and diffusion models to further augment the data-driven capabilities in motion planning. Furthermore, evidence suggests that while imitation learning can enhance motion planning tasks, the imitation metric alone is insufficient as a definitive criterion for assessing the efficacy of motion planners. It is recommended that subsequent research in motion planning incorporate the imitation metric to gauge the intelligence domain of planners and to develop additional metrics for closed-loop evaluation.
@@ A.1 Perception Outputs
At each timestamp, measured states for each tracked entity are provided in terms of positions, velocities, and accelerations. Furthermore, the perception module also provides information about all perceivable landmarks and road boundaries ahead. Stop lines are also essential when the ADV approaches the junction area. The above obstacles and landmarks are either described by coordinates or measured by polynomial curves, referred to as the parametric form. The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its motion status, characterized as non-parametric.
@@ A.2 Scene Rasterization
The state of landmarks and road-bound is rasterized in the same channels, facilitating traffic flow reasoning at intersections. To capture the past motion of all traffic actors, their bounding boxes at consecutive time steps are rasterized on top of map vector layers. Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the fading effect. To distinct ADV from other agents, we rasterize ego’s polygon in an isolated channel, as well as it’s fading history. The rasterized map contains a rendering of traffic light permissibility in junctions, but we tackle it a little bit trick to handle the robustness. we render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static agent’s polygone that exhibit permissibilit of current ego route. All in all, we obtain a 3D tensor of size (Ha , Wa ,C), with C=5 binary channels for the map.
@@ A.3 Instance Vectorization
Parametric information such as surrounding vehicle positions, lane center lines and lane landmarks, are derived with the form of discretized points to depict the traffic scenes. Additionally, route information is obtained to facilitate the ADV in navigating to its intended destination. Given that the vehicle’s historical positions are temporally continuous, whereas the lane center lines, lane landmarks, and route information are spatially continuous, they are structured as vectors $\mathbf{v}$ that connect adjacent points. Specifically, the $i$th vector can be defined as
$\mathbf{v}_{i}=[p_{i},p_{i+1},a_{i}]$ (17)
@@
where $p_{i}$ represents the $i$th point, and $a_{i}$ denotes an attribute that indicates its semantic label such as center lines and visible landmarks.
@@
Similar to the process of scene rasterization, traffic lights are considered as static agents and can be vectorized using unchanged vectors. Each vector can be seen as a node in a graph. Subsequently, a sequence of nodes belonging to the same category will be combined to form an instance, which will then be interconnected within the graph.
A.4 Model structure
@@
Our model architecture consists of two main branches. The first rasterization branch, utilizes four convolutional blocks with varying scales to extract short and long-range features, which are then used to derive the scene embedding. Subsequently, the output from the rasterization branch is fed into two identical-structured decoders, each containing four sub-blocks of deconvolution layers, allowing for step-by-step upsampling to match the size of the input from the rasterization branch. The decoders generate cost volumes and prediction grid maps.
@@
A.5 Route
As indicated previously, the route plays a critical role in guiding ADVs along roadways. Our route information comes from two primary channels: the lane center line and historical commuting trajectories.
The lane center lines, which can be accessed in perception channels, are mainly utilized for cruising scenarios, as the ADV consistently travels along the middle of the lane, and if it deviates from the center, it will be guided back to avoid hazards from adjacent lanes.
However, in lane-changing scenarios, vehicles move across lane markers, making the lane center line ineffective in determining optimal positions for lane changes. Consequently, historical commuting trajectories are utilized to capture the lane-change starting and ending positions. Upon reaching the lane-change starting position or during the lane-change process, the collected trajectory supersedes the center line as the route, providing guidance for the ADV to change lanes.
Likewise, in turning scenarios, the absence of landmarks without HD-maps results in the ADV lacking appropriate instructions to navigate intersections. To address this, the starting and ending positions for turning are recorded. As the ADV approaches intersections, human-driving trajectories are employed as the route to guide the ADV through the intersections.
A.6 Curve Based Sampler
In this appendix, we present the sampling procedure with more details. We follow [48] with little modifications for comfort concerns. Vehicle trajectories can be categorized into two primary types: straight and turning. And turning trajectories encompass lane changes and turning at intersections. Lane-changing trajectories are characterized by variable curvature over time, while turning at intersections follows circular-like curves. As straight lines can be treated as a special case of curves, we introduce the curve-based sampler that focuses on generating straight, constant curvature circle-like, and variable curvature clothoid-like trajectories.
@@ A.6 Curve Based Sampler
Where $\mathbf{s(\xi)}$ is the clothoid curve, $\mathbf{T_{0}}$ and $\mathbf{N_{0}}$ are the tangent and normal vector of point $\mathbf{s_{0}}$, $S(\xi)$ and $C(\xi)$ are associated coefficients and defined as the Fresnel integral. As the shapes of clothoid curves solely determined by the scaling factor a, we can generate a variety of clothoid trajectories by considering a suitable range for a, in combination with the initial curvature obtained from the bicycle model and the speed profiles.
A.7 Lattice Sampler
To begin, this method involves sampling multiple candidate trajectories based on the Frenet Frame, where the smooth lane central line is projected to the S coordinate and its normal direction is the L coordinate. Trajectories are composed of longitudinal and lateral paths, with the longitudinal one described as the station and time, denoted as $S=\left\{s(t)\right\}$, and the lateral path described as the lateral station and longitudinal one denoted, denoted $L={l(s)}$.
Regarding following and overtaking maneuvers, these actions often arise when a preceding vehicle is moving at a slower pace or a new vehicle unexpectedly merges into a lane. The execution of these complex maneuvers is facilitated through the use of an s-t graph, which depicts the occupation of the lane by vehicles over time as illustrated in Figure x. By analyzing the s-t graph, end states can be sampled within unoccupied areas, marked in red (for overtaking) and blue (for following), thus facilitating the completion of these maneuvers through the use of a quintic polynomial.
A.8 Retrieval-based Sampler
@@
To build the expert trajectory dataset, we utilized about 350 hours of driving trajectories from experienced human drivers. However, a substantial portion of the original human driving trajectories contains duplicate or very similar samples. Drawing inspiration from [3], we conducted subdivision and clustering on the original human driving trajectories in order to preserve diversity while simplifying the expert trajectory dataset.
@@
A.9 Hardware

### v3

@@ 4.4 Closed-loop Experiments
Initially, we conducted an ablation study to assess the necessity of various samplers within a closed road environment. Subsequently, we identified the optimal sampler combination for further evaluation on open urban roadways. Comprehensive real-world testing of these methodologies was conducted in densely populated urban settings, under the vigilant supervision of human safety drivers. The specific driving route utilized for our experiments is depicted in Sec. A.15. Throughout the extensive public road testing, spanning over 30 kilometers, the model adeptly executed a diverse array of complex maneuvers, such as lane-following, merging, yielding to pedestrians. The results of these road tests are documented in Tab. 4.
Method comfort event in 30Km efficiency MPI Lat. accel. Jerk $\Delta$Steering Angle Collisions Lost control Discomfort Deviation Minute UniAD[23] 0.351 3.12 3.68 - - - - 15.20 0.31 ST-P3[22] 0.857 4.54 8.53 4 3 14 3 8.62 1.36 GAD-multi 0.489 1.22 2.89 4 1 2 4 5.60 2.72 GAD-lattice 0.728 2.26 4.22 2 0 4 2 5.68 3.75 GAD-base 0.531 1.20 4.21 3 1 3 1 5.56 3.75 GAD-gan 0.511 1.15 3.52 3 1 2 1 5.51 4.28 GAD-gan† 0.492 1.16 2.77 0 1 2 1 5.54 7.50 DPQP 0.284 0.85 2.11 1 0 1 1 5.42 10.0
Table 4: Road test data were systematically collected during a 30 km road test, encompassing the computation of average time per loop. Concurrently, measurements of lateral acceleration, jerk, and steering angle were meticulously recorded. (†) indicates boosted by safety layer and cross frame align. The UniAD consistently deviates and collides with obstacles, thereby failing to accumulate sufficient statistical data.
@@
It is noteworthy that methodologies exhibiting superior imitation metrics do not invariably demonstrate enhanced performance in closed-loop evaluations. This discrepancy may be attributable to the limitations of the imitation metric in adequately assessing the complexities of planning tasks(we illustrated the detail in Sec. A.12). Nonetheless, the multi-modal imitation generator and GAN-based generator significantly enhance the system’s performance in both open-loop and closed-loop settings, primarily in terms of the planner’s intelligence and the continuity between consecutive frames. The efficiency and comfort metrics both show improvements with the integration of data-driven generators. Closed-loop results suggest that multimodality mitigates the brittleness inherent in imitation learning, while the lattice sampler further enhances the robustness of the entire system. See Sec. A.7
@@
To rigorously evaluate the efficacy of the proposed methodology as a substitute for off-the-shelf planners, we conducted a comparative analysis between the GAD and a conventional modularized, user-level, rule-based system, referred to as DPQP in Tab. 4. Notwithstanding, a discernible performance discrepancy persists in terms of system’s robustness and comfortableness. We posit that increase training data size could ameliorate this disparity. The system obtains compariable results with DPQP after boosted by the safety layer and cross frame align as mentioned in Sec. 3.7.
@@ 5 Conclusions
We introduce a sophisticated DL pipeline from prediction to motion planning in an HD map-free setting, augmented with enhanced safety assurance, which has shown substantial enhancements over prevailing data-driven methodologies in Shanghai’s complex urban environments. This paradigm enables the safe deployment of learned planners in real-world scenarios, offers interpretability, and leverages their capacity to evolve with accumulating data, managing more complex scenarios than traditional rule-based systems. Future work may integrate generative learning techniques such as CVAE[27] and diffusion models[40] to further augment these capabilities. Additionally, our results suggest that while imitation learning can enhance motion planning tasks, the imitation metric alone is insufficient as a definitive criterion for assessing the efficacy of motion planners; thus, it is recommended that subsequent research incorporate this metric to gauge the intelligence domain of systems and develop additional metrics for closed-loop evaluation.
@@ A.1 Perception Outputs
At each timestamp, we have 3D detection results for surround view camera, these inputs are fed into an existing tracking system, outputting stat estimates S for all surrounding actors (state encompasses the bounding box dimensions, spatial coordinates, linear velocity, linear acceleration, directional orientation, and the rate of change of heading). Denote a set of discrete times at which tracker outputs state estimates as $[t_{1},t_{2},...t_{T}]$, where time gap between consecutive time steps is constant (the gap is equal to 0.1s for tracker running at a frequency of 10Hz). Furthermore, the perception module also provides information about all perceivable landmarks and road boundaries ahead, measured by the polynomial curve. Stop lines are also essential when the ADV approaches the junction area. The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its motion status(static or movable).
We elucidate the rationale behind our architectural that segregates perception from the comprehensive end-to-end stack: Firstly, for mass production friendly concern, the modularized perception provides explaninablity and tunablitiy. Secondly, information articulated in parametric form facilitates reasoning at the agent level, not solely for application within the safety layer and the analysis of explainable cases, but also to enhance our trajectory generation processes. Thirdly, parameterization encompasses post-tracking and fusion for the results of single or multi-modal detection on raw sensor data which remain formidable unresolved issues, particularly when considering the substitution of these processes with a model to attain comparable results. We leave perception involved version to follow up works when those issue resolved in the future.
@@ A.2 Scene Rasterization
The primary function of scene rasterization is to map scene information to different channels, enabling the neural network to extract scene semantic information more effectively. For example, obstacles, routes, and lane lines are mapped into the RGB channels, respectively. Specifically, the states of landmarks and road-boundary are rasterized in the same channel, facilitating traffic flow reasoning at intersections. To capture the past motions of all traffic actors, their bounding boxes at consecutive time steps are rasterized on top of map vector layers. Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the diminishing visual impact over time. Brightness level at $t$ is equal to $\max[0,(T-t)*\alpha]$, $T$ is timestamp for current frame, where we set $\alpha=0.05$ (the example shown in Fig. 2). To distinct ADV from other agents, we rasterize ego’s polygon in an isolated channel, as well as its fading history. The non-moving obstacles, together with occupancy grids of static label, are rasterized in a different channel since they share similar semantic meaning. The rasterized map contains a rendering of traffic light permissibility in junctions, but we tackle it a little bit trick to handle the robustness. We render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static polygon to exhibit the permissibility of current ego route. The route information is depicted in an isolated channel illustrated in Sec. A.6. All in all, we obtain a 3D tensor of size $(H,W,C)$, with $H=500,W=500,C=5$ for the map.
@@ A.3 Instance Vectorization
Parametric information for certain instance such as trajectories, lanes, and routes are presented in the form of discrete points. In order to enable long-distance interaction, we process these points using a vectorization method to generate a set of instances. For each instance $\mathbf{I}_{i}$, it is comprised of a collection of vectors that belong to the same group:
$\mathbf{I}_{i}=[\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{n}]$ (16)
As for each vector ${\mathbf{v}_{\textit{i}}}$, it can be defined as:
$\mathbf{v}_{i}=[\mathbf{p}_{i},\mathbf{p}_{i+1},a_{i}]$ (17)
@@
where $\mathbf{p}_{i}$ represents the $i$th point coordinates, and $a_{i}$ denotes an attribute that indicates its semantic label such as center lines and visible landmarks.
@@
Similar to the process of scene rasterization, stop lines are considered as static agents and can be vectorized using unchanged vectors. Each vector can be seen as a node in a graph. Subsequently, a sequence of nodes belonging to the same category will be combined to form an instance, which will then be interconnected within the graph.
A.4 Coordinate
The above mentioned representations are built upon Ego-ENU coordinate system, rather than the fixed-oriented coordinate system[9], which is derived by rotating the Ego-ENU coordinate by ADV’s heading. The Ego-ENU coordinate system is centered around the ADV with the x-axis pointing to the east, the y-axis pointing to the north, and the z-axis pointing upwards. In the Ego-ENU coordinate system, the ADV’s historical and future trajectories can be diversely distributed depending on the heading directions, with both x and y coordinates of the same order. This allows for a more balanced representation of the vehicle’s motions in both longitudinal and lateral directions.
Moreover, the fixed-oriented coordinate system tends to extend along the x direction while fluctuating in a small range around 0 along the y direction, leading to challenges in accurately modeling the lateral behaviour such as overtaking and nudging. The scale differences between x and y directions causes lateral errors disproportionately large in comparison to longitudinal behaviours.
Here we represent our comparative experiments to illustrate it. We have conducted three experiments to test performances in Ego-ENU and the fixed-oriented coordinate system respectively. To tackle the unbalanced distribution of trajectories in fixed-oriented one, we present the original and one with feature normalization.
Coordiante system with normalization L2(m)
Ego-ENU x 0.39
Fixed-Oriented x 0.95
Fixed-Oriented $\surd$ 0.63
Table 5: Comparative experiments of coordinate systems.
According to the test results, it is evident that the L2 distance error of the Ego-ENU coordinate system is the lowest since the predicted future trajectory (noted in blue) approximates to the ground truth(noted in white) and closely follows the route(noted in green). Meanwhile, the trajectory in the fixed-oriented coordinate system with feature normalization, also exhibits proximity to the route. However, its longitudinal points are irregularly and non-uniformly distributed. This irregularity is attributed to the scaling of the longitudinal features to the similar order as the lateral ones. Conversely, the trajectory in the fixed-oriented coordinate system without any form of normalization demonstrates the poorest performance, as the end of the trajectory fails to align with the route and deviates outwards, as shown in Fig. 4.
Figure 4: Visualization of comparative results. (a)the predicted trajectory in Ego-ENU closely follows the route. (b)trajectory in the fixed-oriented coordinate system deviates outwards. (c)trajectory in the fixed-oriented coordinate system with feature normalization exhibits proximity to the route.
Therefore, the more symmetrical Ego-ENU coordinate system, which better represents the ADV trajectories, allows for better understanding of vehicle behaviours and enables more effective trajectory planning and overall autonomous driving performance.
A.5 Model structure
@@
Our model architecture consists of two main branches. The first rasterization branch, utilizes four convolutional blocks with varying scales to extract short and long-range features, which are then used to derive the scene embedding. Subsequently, the output from the rasterization branch is fed into two decoders with similar structures, each containing four sub-blocks of deconvolution layers, allowing for step-by-step upsampling to match the size of the input from the rasterization branch. The decoders generate cost volumes and prediction grid maps.
@@
A.6 Route
As indicated previously, the route plays a critical role in guiding ADVs along roadways. Our route information comes from two primary channels: the lane center line and historical commuting trajectories. As shown in Fig. 5.
Figure 5: Route visualization. (a) lane center line is served as route in cruising scenarios. (b) is the turning scenarios, where historical commuting trajectories are served as route, and occupancy provides road boundary information.
The lane center lines, constructed by landmarks which can be accessed in perception outputs, are mainly utilized for cruising scenarios, as the ADV consistently travels along the middle of the lane. We construct three reference lines base on lane center points from current, left and right landmarks to support lane keeping and changing.
However, in turning scenarios, the absence of landmarks without HD-maps results in the ADV lacking appropriate instructions to navigate intersections. To address this, the starting and ending positions from commuting trajectory are recorded. As the ADV approaches intersections, the history human-driving trajectories are employed as the route to guide the ADV through the intersections. Owing to the absence of the road boundary, the scene occupancy has been adopted to delineate the boundary, as illustrated in Fig. 5
Note that in the following visulizations, we omit the occupancy channel for brief.
A.7 Multi-modal Trajectory Generation
We provide qualitive result of single modal, multi-modal and GAD-based generator in Fig. 6 and Fig. 7. GAN-based generator help boosting the multi modality in trajectory generation.
Figure 6: Lane changing scenarios with multi trajectory generation. (a)single trajectory; (b)multi trajectories imitation; (3)GAN-based generator
Figure 7: Right turning scenarios with multi trajectory generation. (a)single trajectory; (b)multi trajectories imitation; (3)GAN-based generator
A.8 Curve-based Sampler
We follow [51] with little modifications for comfort concerns. Vehicle trajectories can be categorized into two primary types: straight and turning. Turning trajectories encompass lane changes and turning at intersections. Lane-changing trajectories are characterized by variable curvatures over time, while turning at intersections follows circular-like curves. As straight lines can be treated as a special case of curves, we introduce the curve-based sampler that focuses on generating straight, constant curvature circle-like, and variable curvature clothoid-like trajectories.
@@ A.8 Curve-based Sampler
Where $\mathbf{s(\xi)}$ is the clothoid curve, $\mathbf{T_{0}}$ and $\mathbf{N_{0}}$ are the tangent and normal vector of point $\mathbf{s_{0}}$, $S(\xi)$ and $C(\xi)$ are associated coefficients and defined as the Fresnel integral. Given the initial curvature obtained from the bicycle model and the speed profiles, we can generate a variety of clothoid trajectories by considering a suitable range for the scaling factor $a$.
A.9 Retrieval-based Sampler
@@
To build the expert trajectory dataset, we utilized about 200,000 driving trajectories from experienced human drivers. However, a substantial portion of the original human driving trajectories contains duplicate or very similar samples. Drawing inspiration from [3], we conducted subdivision and clustering on the original human driving trajectories in order to preserve diversity while simplifying the expert trajectory dataset.
@@
A.10 Lattice Sampler
To begin, this method involves sampling multiple candidate trajectories based on the Frenet frame, where the smooth lane central line is projected to the S coordinate and its normal direction to the L coordinate. Trajectories are composed of longitudinal and lateral paths, with the longitudinal one described as the station and time, denoted as $S=\left\{s(t)\right\}$, and the lateral path described as the lateral station and longitudinal station, denoted as $L={l(s)}$.
Figure 8: Sampling in the Frenet frame, and then converted back to the Cartesian coordinate system.
Regarding following and overtaking maneuvers, these actions often arise when a preceding vehicle is moving at a slower pace or a new vehicle unexpectedly merges into a lane. The execution of these complex maneuvers is facilitated through the use of an s-t graph, which depicts the occupation of the lane by vehicles over time as illustrated in Fig. 9. By analyzing the s-t graph, end states can be sampled within unoccupied areas, marked in red (for overtaking) and blue (for following), thus facilitating the completion of these maneuvers through the use of a quintic polynomial.
Figure 9: ST graphs. (a) An obstacle prepares to cut in the current lane, so ADV chooses to overtake; (b) A preceding vehicle moves slowly, so the ADV slow down.
A.11 Deduction of GAN-based Generator
Given the optimization target for evaluator:
The target for generator become:
$L_{generator}=E(\hat{s})-E(\tilde{s})+Diff(\tilde{s},\hat{s})$ (19)
$E(\tilde{s})$ is determinate in the generator’s step, thus omitted, and we expand $Diff(\tilde{s},\hat{s})$ as defined in Eq. 6 , Eq. 19 become:
$L_{generator}=E(\hat{s})+L2(\tilde{s},\hat{s})+O(\hat{s})+G_{rid}(\hat{s})$ (20)
Where:
$E(\hat{s})=\sum C(\hat{s})+\alpha O_{cc}(\hat{s})+\beta G_{rid}(\hat{s})$ (21)
We then merge $O(\hat{s})$ and $G_{rid}(\hat{s})$ into Eq. 21, tuned by hyper-parameter $\alpha$ and $\beta$, thus Eq. 20 become:
$L_{generator}=E(\hat{s})+L2(\tilde{s},\hat{s})$ (22)
Here the $L2$ is replaced by multi-path imitation defined in Eq. 10 to ensure diversity, the $\hat{s}$ here is given by the generator in Eq. 15, thereby:
$L_{generator}=E(G(m))+L_{imit\_plan}(\tilde{s},G(m))$ (23)
A.12 Limitation of Open-loop Evaluation
We illustrate the limitation of open-loop evaluation. Firstly, a closed-loop planner will adjust to the target in the loop to keep safe and comfortable since real-world planning needs handle the deviations brought by the external disturbances and controller’s modeling errors, which is invisible and omitted, thus cannot be tested by the open-loop planner. Secondly, the validity of the trajectory cannot be verified from the perspective of the downstream controller. The open-loop tests will come to an end once the trajectories are generated and compared with the recorded ones. However, unexpected and unreasonable jerks and curvatures, which can only be tested when the controller is in the loop, are usually ignored; therefore, the results become less convincing. Existing imitation-based approaches formulate the modeling in planning task as prediction problem, which takes ego motion and scene feature to predict future movement. Predominantly, these models are assessed using imitation metrics and subjected to open-loop testing. Nevertheless, a fundamental discrepancy remains, as the aim of planning transcends mere imitation. It is imperative to meticulously craft the learning objectives to avert the entrapment of the model within a local optimum that, while demonstrating proficient motion prediction capacities, fails to adequately fulfill the planning objectives.
A.13 Hardware